OpenStudy (anonymous):

Find the derivative of f(x) = negative 8 divided by x at x = 11.

OpenStudy (mrnood):

\[f(x) = -8x ^{-1}\] that is another way of writing the equation use the standard rules for f'x

OpenStudy (anonymous):

THank you!

OpenStudy (anonymous):

so would it be 8/11

OpenStudy (mrnood):

no - you need to write the equation for f'(x) first, THEN put x=11 into that remember - multiply by the power of x and then reduce the power by 1 (e.g. f(x) = x^2 =>> multiply by 2 and reuce power by one gives f'(x) = 2x) SO multiply by -1 and reduce -1 by 1.....

OpenStudy (mrnood):

-8x^-1 multiply by -1 ==???

OpenStudy (anonymous):

8x?

OpenStudy (mrnood):

no - I am giving you the fisrst step only for now -1*(-8x^-1) = ?

OpenStudy (mrnood):

do oyu know for instance how to do derivative of f(x) = 3x^3 (This is just an example to find out what oyu can already do....) multiply by th eexponent :==> 9x^3 reduce the exponent by 1 ==> 9x^2 so f'(x) = 9x^2 so follow the first step as I showed above: -1*(-8x^-1) = ?

OpenStudy (anonymous):

8/x

OpenStudy (mrnood):

ok now the second step REDUCE the exponent by 1 (8x^-1)